write the equation of a line perpendicular to y=4x-2 and passing through the points(4,2)
Answer by mananth(16949)

1 y = 4 x -2
Divide by 1
y = 4 x + -2
Compare this equation with y=mx+b
slope m = 4

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1


m= - 1/4 ,point ( 4 , 2 )
Find b by plugging the values of m & the point in
y=mx+b
2=-1 + b
b= 3
m= - 1/4
The required equation is y=- ( 1/ 4)x+3
m.ananth@hotmail.ca